1. Three identical metallic conducting spheres carry the following charges: q.-+5.80 pC.q2-+6.20 and q3-_5.20 μ.C The spheres that carry the charges qi and q2 are brought into contact. Then they are separated. After that, one of those two spheres is brought into contact with the third sphere that carries the charge q3; those two are then separated as well a) What is the final charge on the third sphere?
Answer #1

Given q_{1}=+5.8\mu C

q_{2}=+6.2\mu C

and  q_{3}=-5.2\mu C

When q_{1}\ and \ q_{2} are brought in contact ,their charges equalizes

i.e add the charges and divide by 2

(5.8\mu C+6.2\mu C)/2 =12\mu C/2

=6\mu C

Now both spheres  q_{1}\ and \ q_{2} have a charge of 6\mu C

================================

Follow the same steps for q3

(q_{1}+q_{3})/2=(6\mu C-5.2\mu C)/2

=0.8\mu C/2=0.4\mu C

ANSWER: {\color{Red} 0.4\mu C}

===============